OpenStudy (anonymous):

Write log3 6 as a logarithm of base 2.

OpenStudy (anonymous):

you mean \[\log_{3}6 \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well that means of course \[3^x=6\] does it want a decimal? there is no real good answer here unless u are asking somehing i dont know what you mean you can write it as an equality but it will be a decimal

OpenStudy (anonymous):

You mean change bases from log₃(6), to base 2 ok changing bases that would be equal to. log₃(6)=log₂(6)/log₂(3) Hope you understand

OpenStudy (anonymous):

\[\ln 2/6=x\]

OpenStudy (anonymous):

yea i mean we are both correct here but im not sure exactly what you are asking for